A plane flies 600 miles upwind in 3 hours. It makes the return trip in only 2 hours. What is the speed of the plane still in the air?
Answer by mananth(16949)

Plane speed =x mph
wind speed =y mph
against wind 3 hours
with wind 2 hours

Distance with wind 600 miles distance against wind 600
t=d/r against wind (x-y)
600.00 / ( x - y )= 3.00

3.00 x - -3.00 y = 600.00 ....................1
with wind (x+y)
600.00 / ( x + y )= 2.00
2.00 ( x + y ) = 600.00
2.00 x + 2.00 y = 600.00 ...............2
Multiply (1) by 2.00
Multiply (2) by 3.00
we get y
6.00 x + -6.00 y = 1200.00
6.00 x + 6.00 = 1800.00
12.00 x = 3000.00
/ 12.00
x = 250 mph

plug value of x in (1) y
3 x -3 y = 600
750 -3 -750 = 600
-3 y = 600
-3 y = -150 mph
y = 50
Plane 250 mph
wind 50 mph